A 2kg steel ball hangs on a 8.0-m-long string.  The ball is pulled sideways to an angle of 45°, then released.  At the very bottom of the swing the ball strikes a 3kg steel block (perfectly elastic collision) that is resting on a frictionless table with a spring (k=20,000 N/m) further down.  

(a 10 points) To what maximum angle does the ball rebound (degrees)?

b 5 points) What is the compression in the spring (m)?
Answer #1

To answer the first part we need to calculate the initial height of the steel ball. Then we use energy conservation to find out the velocity of the ball before collision. Then use equations of elastic collision to find rebound velocity and then rebound angle.

In second part use the velocity of the block after collision and then energy conservation.

a Caso LW L W = L - L Coso = L(1-650) Criven L = 8m, d= 45° :h=8C1-C6s 450) = 2.343 m = 1289.8x2.343 Velocity of Skeel ball j